OpenStudy (anonymous):

please help ill give a medal C=5/9(f-32) solve for F

OpenStudy (anonymous):

Nice weather to day. Think about how to convert from Celsius to Fahrenheit

OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F. C = 5 over 9(F - 32) Which of the following formulas correctly solves for F? F = 9 over 5C + 32 F = 9 over 5C - 32 F = 9C + 32 over 5 F = 9C - 32 over 5

OpenStudy (anonymous):

@undeadknight26 if you want to help me here is the link to my other Question